Every dollar saved in purchasing is equivalent to a dollar of new income. Material cost-savings have a greater influence on the ROA than do sales increases, therefore it is easier for companies to generate cost-savings than to increase market share. To prove the aforementioned statements, use the given information of ABC Company to calculate your answers for the following questions. (Show your calculations) (20 points)

Sales $4,900,000                                             Labor Cost $600,000

Fixed Assets $3,000,000                                 Overhead Cost $900,000

Inventories $450,000                                       Materials Cost $2,400,000

Accounts Receivable $275,000                       Other Costs $700,000

Cash $250,000

a.    What is the current profit margin of the company? (3 points)

b.    What is the current investment turnover of the company? (3 points)

c.    What is the current ROA of the company? (3 point)

d.    What will be the new profit margin if purchasing is able to reduce 5% of materials costs? (3 points)

e.    What will be the new investment turnover if purchasing is able to reduce 5% of materials costs? (3 points)

f.     What will be the new ROA if purchasing is able to reduce 5% of materials costs? (3 point)

g.    What is your conclusion to this question? (2 point)
Answer #1
Sales 4900000 Labour 600000
FA 3000000 Overhead 900000
Inventory 450000 Material 2400000
Acc Rec 275000 Other costs 700000
Cash 250000 Total costs 4600000
a) Profit = sales - total costs = 4900000 - 4600000
300000
Hence profit margin = profit/sales
6.1%
b. Investment turnover = sales/total assets
Total assets = Fixed assets + inventory + acc rec + cash
3975000
Hence investment turnover = 4900000/3975000
             1.23
c. ROA or return on assets = total income/total assets
=300000/3975000 7.5%
d. New profit margin if 5% material cost is reduced:
new material cost =2400000*0.95 2280000
i.e reduction by 2400000 - 2280000 120000
Hence total cost = 4600000 - 12000 4480000
Hence profit = 4900000 - 4480000 420000
Profit margin = 420000/4900000 8.6%
Increases from 6.1% to 8.6%
e. Impact on investment turnover if material cost reduces by 5%
Investment turnover is given as sales/total assets
since material cost reduction did not have an impact on either sales or total assets
this ratio remains the same as earlier i.e. 1.23
f. New return on assets = 420000/3975000
10.6%
g. Thus a 5% reduction in material cost has increase the ROA from 7.5% to 10.6%
this increase is tremendous 10.6%/7.5% = 140.0%
i.e improvement in ROA by 40% with a 5% redn in material cost